Southwestern Pipe Supply (SPS) is a distributor of pipes, pipe lining and coating. SPS buys couplings...

  1. Southwestern Pipe Supply (SPS) is a distributor of pipes, pipe lining and coating. SPS buys couplings from Baker Corp. for $12 each. It costs $3000 to store couplings at the warehouse (independent of the number of units stored, similar to rent). A contractor is willing to buy the couplings at $15 each.
    1. How many couplings should SPS sell to the contractor before they breakeven. Show all calculations. Specify units.
    2. How much would be the net income or profit if SPS sold 1350 couplings to the contractor?

On a separate graph paper, provide the graphical representation of the breakeven point. Provide appropriate labels and numerical values for the X and Y axes, cost and revenue lines, and the breakeven point
Answer #1

The breakeven point formula is = Fixed cost / (Selling price – Variable cost)

Profit formula = (Selling price – Variable cost)*Number of units – Fixed cost

Fixed cost = $3000

Selling price = $15

Variable cost = $12

The breakeven point = 3000/(15-12) = 1000 units

In case 1350 units are sold then the profit will be

Profit = (15-12)*1350 – 3000 = $1050

The chart is shown below

Chart 1 : x x cronimnos e as A B C D Ε G H K L M N O P Q R т 200 200 Cost-Revenue - Cost — Revenue 30000 25000 Units Cost Rev